I need help solving this word problem we are to use a system of linear equations to solve the problem.
Your teacher is giving you a test worth 100 points containing 40 questions. There are two-point questions and four-point questions on the test. How many of each type of question are on the test?
Please help this is really important!! thanks

Answer by richwmiller(17219)

We have two unknowns. number of 2 point and 4 point questions.
t +f=40
t=40-f
2t+4f=100
substitute 40-f for t
2*(40-f)+4f=100
f=10 and t=30